# A Generalized Back-Door Criterion for Linear Regression

### v1

@@ Abstract
What assumptions about the data-generating process are required to permit a causal interpretation of partial regression coefficients? To answer this question, this paper generalizes Pearl’s single-door and back-door criteria and proposes a new criterion, which enables the identification of total or partial causal effects. In addition, this paper elucidates the mechanism of post-treatment bias, showing that a repeated sequence of nodes can be a potential source of this bias. The results apply to linear data-generating processes represented by directed acyclic graphs with distribution-free error terms.
@@ 1 Introduction
To analyze causal structures, we need to make certain assumptions about them. For example, Wright1921 introduced path analysis and illustrated it using a path diagram. This method has evolved through several studies, including Haavelmo1943, Joreskog1970, and Goldberger1972, and is also known as linear structural equation modeling (SEM). It is now widely used to analyze linear causal relationships.
As a generalization of this approach, Pearl1995 developed a framework for causal inference that can also be applied to non-linear causal structures. Within this framework, Pearl1998 applied the back-door criterion—originally proposed by Pearl1993 as a method for covariate selection—to linear data-generating processes represented by directed acyclic graphs (DAGs), showing that it provides a sufficient condition under which a regression coefficient coincides with a total effect in the sense of SEM. The back-door criterion has since been generalized by Shpitser2012, Zander2014, Maathuis2015, Perkovic2015, and Perkovic2020. In addition to back-door adjustment, other approaches, such as front-door adjustment (Pearl1995) and conditional instrumental variables (BritoandPearl2002), have been established for identifying total effects.
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. Foygel2012 proposed the half-trek criterion and discussed the identifiability of path coefficients based on the covariance structure. In addition to the studies mentioned above, various other previous works exist; however, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
Note, however, that the scope of this paper is limited to linear causal structures, as the main objective is to derive a broad sufficient condition that permits causal interpretations of partial regression coefficients. Therefore, it remains an open question whether similar results hold for probability distributions of causal models that allow for nonlinearity.
@@ 2.2 Regression on the Population
is called the population partial regression coefficient in the regression of $Y$ on $\bm{X}$.
@@ 3.2 Main Results
Condition $(iii)$ is particularly easy to check, because $\bm{S}\setminus\{X_{j}\}$ blocks a back-door path from $X_{j}$ to $X_{i}$ that contains no v-structure if and only if $X_{i}\notin\bm{S}$ and at least one node $X_{k}\in\bm{S}\setminus\{X_{j}\}$ lies on that back-door path.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on variables affected by a treatment (i.e., intervened variables) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the quasi–back–door criterion is required.
@@
Although bias is often defined as the difference from the total effect $\tau_{ij}$, this paper adopts the above definition because it is consistent with the interpretation of partial regression coefficients introduced at the beginning of the Introduction, which is the main focus of the present study.
@@ 4 Examples
Another bias $\gamma_{41|3}$ can be explained by two sequences of nodes: $X_{1}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$ and $X_{1}\rightarrow X_{2}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$. The former has long been recognized as a source of distortion in causal inference, as captured by the concept of path blocking. In contrast, the latter corresponds to the structure clarified in this paper and has not been widely discussed in the literature. In this model, the bias arises except when $a_{31}=-a_{32}a_{21}$.
@@ 5 Proofs
Even if $X_{s}$ is a collider, its descendant $X^{\prime}_{p}$ belongs to $\bm{S}\setminus\{X^{\prime}_{q}\}$. Hence, these back-door paths remain unblocked by $\bm{S}\setminus\{X^{\prime}_{q}\}$, contradicting assumption $(ii)$. Therefore, conditions (a) and (b) cannot both hold.
@@
where $\gamma_{i0}$ and $\gamma_{im}$ are the population partial regression coefficients obtained from regressing $e_{i}$ on the random vector consisting of the constant $1$ and all elements of $\bm{Z}$. The existence of $\gamma_{i0}$ and $\gamma_{im}$ is guaranteed by Proposition 2.6, and $\mathop{\rm Cov}\nolimits(X_{m},e^{*}_{i})=0$ follows from Proposition 2.4. Here, the notation $\gamma_{im}$ is consistent with that used for the bias terms in Theorem 3.5, since they correspond to the same quantities.

### v2

@@ Abstract
What assumptions about the data-generating process are required to permit a causal interpretation of partial regression coefficients? To answer this question, this paper generalizes Pearl’s single-door and back-door criteria and proposes a new criterion that enables the identification of total or partial causal effects. In addition, this paper elucidates the mechanism of post-treatment bias, showing that a repeated sequence of nodes can be a potential source of this bias. The results apply to linear data-generating processes represented by directed acyclic graphs with distribution-free error terms.
@@ 1 Introduction
To analyze causal structures, we need to make certain assumptions about them. For example, Wright1921 introduced path analysis and illustrated the structures using path diagrams. This method has evolved through several studies, including Haavelmo1943, Joreskog1970, and Goldberger1972, and is also known as linear structural equation modeling (SEM). It is now widely used to analyze linear causal relationships.
As a generalization of this approach, Pearl1995 developed a framework for causal inference that can also be applied to non-linear causal structures. Within this framework, Pearl1998 applied the back-door criterion—originally proposed by Pearl1993 as a method for covariate selection—to linear data-generating processes represented by directed acyclic graphs (DAGs), showing that it provides a sufficient condition under which a regression coefficient coincides with a total effect in the sense of SEM. The back-door criterion has since been generalized by Shpitser2012, Zander2014, Maathuis2015, Perkovic2015, and Perkovic2020. In addition to the back-door criterion, other approaches, such as the front-door criterion (Pearl1995) and conditional instrumental variables (BritoandPearl2002), have been established for identifying total effects.
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. Tian2012, Foygel2012, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. In addition to the studies mentioned above, various other previous works exist; however, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
Note, however, that the scope of this paper is limited to linear causal structures, as the main objective is to derive a broad sufficient condition that permits causal interpretations of partial regression coefficients. Therefore, it remains an open question whether similar results hold for the probability distributions of causal models that allow for nonlinearity.
@@ 2.2 Regression on the Population
is referred to as the population partial regression coefficient in the regression of $Y$ on $\bm{X}$.
@@ 3.2 Main Results
Condition $(iii)$ is particularly easy to check because $\bm{S}\setminus\{X_{j}\}$ blocks a back-door path from $X_{j}$ to $X_{i}$ that contains no v-structure if and only if $X_{i}\notin\bm{S}$ and at least one node $X_{k}\in\bm{S}\setminus\{X_{j}\}$ lies on that back-door path.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on variables affected by a treatment (i.e., intervened variables) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
@@
Although bias is often defined as the difference from the total effect $\tau_{ij}$, this paper adopts the above definition because it is consistent with the interpretation of partial regression coefficients introduced at the beginning of Section 1, which is the main focus of the present study.
@@ 4 Examples
Another bias $\gamma_{41|3}$ can be explained by two sequences of nodes: $X_{1}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$ and $X_{1}\rightarrow X_{2}\rightarrow X_{3}\leftarrow X_{2}\rightarrow X_{4}$. The former has long been recognized as a source of distortion in causal inference, as captured by the concept of path blocking. In contrast, the latter corresponds to the structure clarified in this paper and has not been widely discussed in the literature. In this model, the bias arises except when $a_{31}=-a_{32}a_{21}$. Even when $X_{2}$ is not observable, fortunately, the total effect $\tau_{31}=a_{31}+a_{32}a_{21}$ can be identified by the simple regression coefficient of $X_{1}$ on $X_{3}$, which allows for a certain degree of consideration of the influence of $\gamma_{41|3}$.
@@ 5 Proofs
Even when $X_{s}$ is a collider, its descendant $X^{\prime}_{p}$ belongs to $\bm{S}\setminus\{X^{\prime}_{q}\}$. Hence, these back-door paths remain unblocked by $\bm{S}\setminus\{X^{\prime}_{q}\}$, contradicting assumption $(ii)$. Therefore, conditions (a) and (b) cannot both hold.
@@
where $\gamma_{i0}$ and $\gamma_{im}$ are the population partial regression coefficients obtained from regressing $e_{i}$ on the random vector consisting of the constant $1$ and all elements of $\bm{Z}^{*}$. The existence of $\gamma_{i0}$ and $\gamma_{im}$ is guaranteed by Proposition 2.6, and $\mathop{\rm Cov}\nolimits(X_{m},e^{*}_{i})=0$ follows from Proposition 2.4. Here, the notation $\gamma_{im}$ is consistent with that used for the bias terms in Theorem 3.5, since they correspond to the same quantities.